# New results on $k$-Roudenff’s conjecture

### v1

@@ 1. Introduction
In this paper we study the number of $k$-neighborly reorientations of a rank $r$ oriented matroid $\mathcal{M}$ on $n$ elements, denoted by $f_{\mathcal{M}}(r,n,k)$. We obtain some partial results on $k$-Roudneff’s conjecture (proposed in [21]), the case $k=1$ being a well-known conjecture stated originally in terms of projective pseudohyperplane arrangements [33].
Let us briefly give some basic notions and definitions on oriented matroid theory needed for the rest of the paper. We refer the reader to [3] for background on oriented matroid theory. Given an oriented matroid $\mathcal{M}=(E,\mathcal{C})$ of a finite ground set $E$ and a set of sign-vectors $\mathcal{C}\subseteq\{+,-,0\}^{E}$ called circuits, the size of a member $X\in\{+,-,0\}^{E}$ is the size of its support $\underline{X}=\{e\in E\mid X_{e}\neq 0\}$. Throughout the paper all oriented matroids are considered simple, i.e., all circuits have size at least $3$. The rank $r$ of $\mathcal{M}$ is the size of a largest set not containing a circuit of $\mathcal{M}$. An oriented matroid of rank $r$ is called uniform if all its circuits are of size $r+1$. For a sign-vector $X\in\{+,-,0\}^{E}$ on ground set $E$ we denote by $X^{+}:=\{e\in E\mid X_{e}=+\}$ the set of positive elements of $X$ and $X^{-}:=\{e\in E\mid X_{e}=-\}$ its set of negative elements. Hence, the set $\underline{X}=X^{+}\cup X^{-}$ is the support of $X$. For a subset $R\subseteq E$ the reorientation of $R$ is the oriented matroid $_{-R}\mathcal{M}$ obtained from $\mathcal{M}$ by reversing the sign of $X_{e}$ for every $e\in R$ and $X\in\mathcal{C}$. The set of all oriented matroids that can be obtained this way from $\mathcal{M}$ is the reorientation class $\mathcal{M}$] of $\mathcal{M}$. We denote by X-X the sign-vector $_{-E}X$ where all signs are reversed, i.e., such that $X^{+}=X^{-}$ and $X^{-}=X^{+}$. We say that $X$ is positive if $X^{-}=\emptyset$ and $\underline{X}\neq\emptyset$.
@@
An oriented matroid $\mathcal{M}=(E,\mathcal{C})$ is acyclic if every circuit has positive and negative signs, i.e., if $|X^{+}|>0$ and $|X^{-}|>0$ for every circuit $X\in\mathcal{C}$. Given an integer $k\geq 0$, we say that $\mathcal{M}$ is $k$-neighborly if $|X^{+}|>k$ and $|X^{-}|>k$ for every $X\in\mathcal{C}$, in other words, if for every subset $S\subseteq E$ of size at most $k$ the reorientation $_{-S}\mathcal{M}$ is acyclic. A reorientation $R$ of $\mathcal{M}$ is a $k$-neighborly reorientation if $_{-R}\mathcal{M}$ is $k$-neighborly. It follows from the definition that a $k$-neighborly oriented matroid is $k^{\prime}$-neighborly for all $0\leq k^{\prime}\leq k$. Note that $\mathcal{M}$ is 0-neighborly if and only if $\mathcal{M}$ is acyclic. If $\mathcal{M}$ is $1$-neighborly, then $\mathcal{M}$ is called matroid polytope. If $\mathcal{M}$ is a rank $r$, $k$-neighborly oriented matroid, then $r\geq 2k+1$. If $r=2k+1$ then $\mathcal{M}$ is often just called neighborly (i.e., when $k$ is the maximum possible). There is quite some work about neighborly oriented matroids, starting with Sturmfels [34] and [3, Section 9.4] but also more recent works such as [28, 30]. In the realizable setting, a $k$-neighborly matroid $\mathcal{M}$ is a $k$-neighborly polytope, important mathematical objects of study in discrete geometry (i.e., a polytope such that every set of at most $k$ vertices are the vertices of a face). Also in the realizable setting, a $k$-neighborly reorientation of $\mathcal{M}$ corresponds to a permissible projective transformation that sends $\mathcal{M}$ to a $k$-neighborly polytope (see [6, Theorem 1.2] and [14]).
@@
The cyclic polytope of dimension $d$ with $n$ vertices discovered by Carathéodory [4], is the convex hull in $\mathbb{R}^{d}$ of $n\geq d+1\geq 3$ different points $x(t_{1}),\dots,x(t_{n})$ of the moment curve $\mu:\mathbb{R}\longrightarrow\mathbb{R}^{d},\ t\mapsto(t,t^{2},\dots,t^{d})$. Cyclic polytopes are neighborly polytopes and play an important role in the combinatorial convex geometry due to their connection with certain extremal problems, for example, the upper bound theorem due to McMullen [27]. The oriented matroid associated to the cyclic polytope of dimension d=r1d=r-1 with $n$ elements is the alternating oriented matroid ${C}_{r}(n)$. It is the realizable uniform oriented matroid of rank $r$ and ground set $E=[n]:=\{1,\ldots,n\}$ such that every circuit $X\in\mathcal{C}$ is alternating, i.e., $X_{i_{j}}=-X_{i_{j+1}}$ for all $1\leq j\leq r$ if $\underline{X}=\{i_{1},\ldots,i_{r+1}\}$ and $i_{1}<\ldots<i_{r+1}$.
Let denote $c_{r}(n,k)=f_{{C}_{r}(n)}(r,n,k)$, i.e., $c_{r}(n,k)$ is the number of $k$-neighborly reorientations of the rank r alternating oriented matroid on $n$ elements. In 1991 Roudneff [33] proved that $c_{r}(n,1)\geq{2}\sum_{i=0}^{r-3}\binom{n-1}{i}$ and that $c_{r}(n,1)={2}\sum_{i=0}^{r-3}\binom{n-1}{i}$ if $n\geq 2r-1$. Recently, it was computed the exact number of $c_{r}(n,k)$ for $n\geq 2(r-k)+1$ and $r\geq 2k+1\geq 3$[21, Theorem 3.11]:
@@
In fact, in [21], Theorem 3.11 proved that $o(C_{r}(n),i)=2\binom{n}{r-1-2i}$ for every $i=k,\ldots,\lfloor\frac{r-1}{2}\rfloor$ and $n\geq 2(r-k)+1$, where $o(C_{r}(n),i)$ was defined as the the number of $i$-neighborly reorientations of $C_{r}(n)$ that are not $(i+1)$-neighborly. So, for $n\geq 2(r-k)+1$ we have that $c_{r}(n,k)=2\sum\limits_{i=k}^{\lfloor\frac{r-1}{2}\rfloor}\binom{n}{r-1-2i}=2\sum\limits_{i=0}^{r-1-2k}\binom{n-1}{i}$ as mentioned above. We then notice that the value of $c_{r}(n,k)$ is unknown for $n\leq 2(r-k)$.
@@ 1.2. $k$-Roudneff’s conjecture
Roudneff conjectured that the rank $r$ alternating oriented matroid on $n\geq 2r-1$ elements has the maximum number of $1$-neighborly reorientations over all rank $r$ oriented matroids on $n$ elements [33, Conjecture 2.2]. Here we point out that the conjecture was stated originally in terms of projective pseudohyperplane arrangements and we state here in terms of oriented matroids. Many of the combinatorial properties of projective pseudohyperplane arrangements can be studied in the language of oriented matroids. Indeed, an oriented matroid on $n$ elements of rank $r$ is naturally associated with some arrangement of $n$ (pseudo) hyperplanes in the projective space of dimension d=r1d=r-1, and conversely the (simple) oriented matroids are precisely those associated with some arrangement of pseudo hyperplanes [11]. We refer the reader to [21] for more details on the translation of Roudneff’s conjecture from arrangements to oriented matroids.
@@
The conjecture remains open for $r\geq 6$. The case $r=3$ is not difficult to prove. Ramírez Alfonsín proved the case $r=4$ in 1999 [31] and recently Hernández-Ortiz et al. proved it for $r=5$ [20]. In 2015 Montejano and Ramírez Alfonsín proved the conjecture for Lawrence oriented matroids [29]. Furthermore, Bárány et al.[2] proved that for fixed $r$, the number of $1$-neighborly reorientations of any rank $r$ realizable oriented matroid on $n$ elements is $2\binom{n-1}{r-3}+2O(n^{r-4})$, i.e., Roudneff’s conjecture holds asymptotically in the realizable setting (here we also point out that the authors proved this result in terms of hyperplane arrangements in the Euclidean space and not in terms of realizable oriented matroids). In this manuscript, we prove that Roudneff’s conjecture holds asymptotically for all oriented matroids (Section 3, Corollary 3.10).
Conjecture 1.1 was stated for $n\geq 2r-1$ probably because in 1991 the value of $c_{r}(n,1)$ was unknown for n<2r1n<2r-1 but in 2001 was computed its value for $n\geq r+1$ [13]. So, in view of Roudneff’s conjecture, the authors asked in [29] if Roudneff’s conjectured holds for $n\geq r+1$ [29, Question 2] and again, it turns out that it is true for $r\leq 5$ [20] and for Lawrence oriented matroids [29]. For an integer $k\geq 0$ it was proposed in [21] the following strengthening of 1.1.
@@
Firstly, notice that $f_{\mathcal{M}}(r,n,k)=f_{\mathcal{M}^{\prime}}(r,n,k)$ if $\mathcal{M}^{\prime}\in[\mathcal{M}]$. So, it is enough to check Conjecture 1.2 for the reorientation classes $\mathcal{M}$] of $\mathcal{M}$. Secondly, every rank $r$ oriented matroid $\mathcal{M}$ on $n$ elements can be perturbed to become a uniform rank $r$ oriented matroid $\mathcal{M}^{\prime}$ on $n$ elements, see [3, Corollary 7.7.9] and in particular $f_{\mathcal{M}}(r,n,k)\leq f_{\mathcal{M}^{\prime}}(r,n,k)$ (see [21, Theorem 3.16]). Thirdly, it is well-known that there is exactly one reorientation class of uniform rank $r$ oriented matroids on $n\leq r+2$ elements (see [21, Remark 1.3] and [3, Section 6.1]). Thus, we obtain the following observation.
@@
For $k=1$ 1.2 combines Roudneff’s conjecture and [29, Question 2]. Hence, for $k=1$ the answer is positive if $r\leq 5$ and also holds for Lawrence oriented matroids. Conjecture 1.2 was also proved in [21] for $k=2$ and $r=6$, and also for every odd rank $r=2k+1$ [21].
One might think that as in the Upper Bound Theorem [27] the reorientation class of $\mathcal{C}_{r}(n)$ is unique in attaining the maximum in 1.2. However, there are different reorientation classes that attain $c_{r}(n,k)$, for $r=5,n=8,9$ and $k=2$ and also for $r=7,n=10$ and $k=3$ [21, Theorem 4.1]. Nevertheless, for $k=1$ it was conjectured in [20] that the reorientation class of ${C}_{r}(n)$ is the unique class with $c_{r}(n,k)$ $k$-neighborly reorientations.
@@ 2. $k$-neighborly reorientations of LOMs
In this section we study the parameter $f_{\mathcal{M}}(r,n,k)$ when $\mathcal{M}$ is a Lawrence oriented matroid (LOM). We obtain obtain some results on $k$-Roudenff’s conjecture for LOMs.
Before defining a LOM, we need to introduce the chirotope of an oriented matroid. The definition of the base axioms of an oriented matroid is due to Gutierrez Novoa [16] who called the structure multiply ordered sets. Lawrence [26] proved that this yields another axiom system for oriented matroids. The term chirotope of an oriented matroid is due to Dress [8], who reinvented oriented matroids. An oriented matroid of rank $r$ is a pair $\mathcal{M}=(E,\chi)$ of a finite ground set $E$ and a chirotope $\chi:E^{r}\to\{+,0,-\}$ satisfying:
@@
If $\chi:E^{r}\rightarrow\{-,+\}$ is a chirotope, then $\mathcal{M}=(E,\chi)$ is a uniform oriented matroid. Moreover, if $E=\{1,\ldots,n\}$ and $\chi(B)=+$ for any ordered tuple $B=(b_{1}<\ldots<b_{r})$, then $\mathcal{M}=(E,\chi)$ is ${C}_{r}(n)$, the alternating oriented matroid of rank $r$ on $n$ elements. It is known that every oriented matroid $\mathcal{M}$ has exactly two basis orientations and these two basis orientations are opposite, $\chi$ and $\chi$ (see [3, Proposition 3.5.2]).

### v2

# New results on $k$-Roudneff’s conjecture

@@ 1. Introduction
In this paper we study the number of $k$-neighborly reorientations of a rank $r$ oriented matroid $\mathcal{M}$ on $n$ elements, denoted by $f_{\mathcal{M}}(r,n,k)$. We obtain some partial results on $k$-Roudneff’s conjecture (proposed in [HKM24]), the case $k=1$ being a well-known conjecture stated originally in terms of projective pseudohyperplane arrangements [R91].
Let us briefly give some basic notions and definitions on oriented matroid theory needed for the rest of the paper. We refer the reader to [BVSWZ99] for background on oriented matroid theory. Given an oriented matroid $\mathcal{M}=(E,\mathcal{C})$ of a finite ground set $E$ and a set of sign-vectors $\mathcal{C}\subseteq\{+,-,0\}^{E}$ called circuits, the size of a member $X\in\{+,-,0\}^{E}$ is the size of its support $\underline{X}=\{e\in E\mid X_{e}\neq 0\}$. Throughout the paper all oriented matroids are considered simple, i.e., all circuits have size at least $3$. The rank $r$ of $\mathcal{M}$ is the size of a largest set not containing a circuit of $\mathcal{M}$. An oriented matroid of rank $r$ is called uniform if all its circuits are of size $r+1$. For a sign-vector $X\in\{+,-,0\}^{E}$ on ground set $E$ we denote by $X^{+}:=\{e\in E\mid X_{e}=+\}$ the set of positive elements of $X$ and $X^{-}:=\{e\in E\mid X_{e}=-\}$ its set of negative elements. Hence, the set $\underline{X}=X^{+}\cup X^{-}$ is the support of $X$. For a subset $R\subseteq E$ the reorientation of $R$ is the oriented matroid $_{-R}\mathcal{M}$ obtained from $\mathcal{M}$ by reversing the sign of $X_{e}$ for every $e\in R$ and $X\in\mathcal{C}$. The set of all oriented matroids that can be obtained this way from $\mathcal{M}$ is the reorientation class $\mathcal{M}$] of $\mathcal{M}$. We denote by X-X the sign-vector $_{-E}X$ where all signs are reversed, i.e., such that $X^{+}=X^{-}$ and $X^{-}=X^{+}$. We say that $X$ is positive if $X^{-}=\emptyset$ and $\underline{X}\neq\emptyset$.
@@
An oriented matroid $\mathcal{M}=(E,\mathcal{C})$ is acyclic if every circuit has positive and negative signs, i.e., if $|X^{+}|>0$ and $|X^{-}|>0$ for every circuit $X\in\mathcal{C}$. Given an integer $k\geq 0$, we say that $\mathcal{M}$ is $k$-neighborly if $|X^{+}|>k$ and $|X^{-}|>k$ for every $X\in\mathcal{C}$, in other words, if for every subset $S\subseteq E$ of size at most $k$ the reorientation $_{-S}\mathcal{M}$ is acyclic. A reorientation $R$ of $\mathcal{M}$ is a $k$-neighborly reorientation if $_{-R}\mathcal{M}$ is $k$-neighborly. It follows from the definition that a $k$-neighborly oriented matroid is $k^{\prime}$-neighborly for all $0\leq k^{\prime}\leq k$. Note that $\mathcal{M}$ is 0-neighborly if and only if $\mathcal{M}$ is acyclic. If $\mathcal{M}$ is $1$-neighborly, then $\mathcal{M}$ is called matroid polytope. If $\mathcal{M}$ is a rank $r$, $k$-neighborly oriented matroid, then $r\geq 2k+1$. If $r=2k+1$ then $\mathcal{M}$ is often just called neighborly (i.e., when $k$ is the maximum possible). There is quite some work about neighborly oriented matroids, starting with Sturmfels [S88] and [BVSWZ99, Section 9.4] but also more recent works such as [MP15, P13]. In the realizable setting, a $k$-neighborly matroid $\mathcal{M}$ is a $k$-neighborly polytope, important mathematical objects of study in discrete geometry (i.e., a polytope such that every set of at most $k$ vertices are the vertices of a face). Also in the realizable setting, a $k$-neighborly reorientation of $\mathcal{M}$ corresponds to a permissible projective transformation that sends $\mathcal{M}$ to a $k$-neighborly polytope (see [CS85, Theorem 1.2] and [GL15]).
@@
The cyclic polytope of dimension $d$ with $n$ vertices discovered by Carathéodory [Car], is the convex hull in $\mathbb{R}^{d}$ of $n\geq d+1\geq 3$ different points $x(t_{1}),\dots,x(t_{n})$ of the moment curve $\mu:\mathbb{R}\longrightarrow\mathbb{R}^{d},\ t\mapsto(t,t^{2},\dots,t^{d})$. Cyclic polytopes are neighborly polytopes and play an important role in the combinatorial convex geometry due to their connection with certain extremal problems, for example, the upper bound theorem due to McMullen [Mac]. The oriented matroid associated to the cyclic polytope of dimension d=r1d=r-1 with $n$ elements is the alternating oriented matroid ${C}_{r}(n)$. It is the realizable uniform oriented matroid of rank $r$ and ground set $E=[n]:=\{1,\ldots,n\}$ such that every circuit $X\in\mathcal{C}$ is alternating, i.e., $X_{i_{j}}=-X_{i_{j+1}}$ for all $1\leq j\leq r$ if $\underline{X}=\{i_{1},\ldots,i_{r+1}\}$ and $i_{1}<\ldots<i_{r+1}$.
Let denote $c_{r}(n,k)=f_{{C}_{r}(n)}(r,n,k)$, i.e., $c_{r}(n,k)$ is the number of $k$-neighborly reorientations of the rank r alternating oriented matroid on $n$ elements. In 1991 Roudneff [R91] proved that $c_{r}(n,1)\geq{2}\sum_{i=0}^{r-3}\binom{n-1}{i}$ and that $c_{r}(n,1)={2}\sum_{i=0}^{r-3}\binom{n-1}{i}$ if $n\geq 2r-1$. Recently, it was computed the exact number of $c_{r}(n,k)$ for $n\geq 2(r-k)+1$ and $r\geq 2k+1\geq 3$[HKM24, Theorem 3.11]:
@@
In fact, in [HKM24], Theorem 3.11 proved that $o(C_{r}(n),i)=2\binom{n}{r-1-2i}$ for every $i=k,\ldots,\lfloor\frac{r-1}{2}\rfloor$ and $n\geq 2(r-k)+1$, where $o(C_{r}(n),i)$ was defined as the the number of $i$-neighborly reorientations of $C_{r}(n)$ that are not $(i+1)$-neighborly. So, for $n\geq 2(r-k)+1$ we have that $c_{r}(n,k)=2\sum\limits_{i=k}^{\lfloor\frac{r-1}{2}\rfloor}\binom{n}{r-1-2i}=2\sum\limits_{i=0}^{r-1-2k}\binom{n-1}{i}$ as mentioned above. We then notice that the value of $c_{r}(n,k)$ is unknown for $n\leq 2(r-k)$.
@@ 1.2. $k$-Roudneff’s conjecture
Roudneff conjectured that the rank $r$ alternating oriented matroid on $n\geq 2r-1$ elements has the maximum number of $1$-neighborly reorientations over all rank $r$ oriented matroids on $n$ elements [R91, Conjecture 2.2]. Here we point out that the conjecture was stated originally in terms of projective pseudohyperplane arrangements and we state here in terms of oriented matroids. Many of the combinatorial properties of projective pseudohyperplane arrangements can be studied in the language of oriented matroids. Indeed, an oriented matroid on $n$ elements of rank $r$ is naturally associated with some arrangement of $n$ (pseudo) hyperplanes in the projective space of dimension d=r1d=r-1, and conversely the (simple) oriented matroids are precisely those associated with some arrangement of pseudo hyperplanes [FL78]. We refer the reader to [HKM24] for more details on the translation of Roudneff’s conjecture from arrangements to oriented matroids.
@@
The conjecture remains open for $r\geq 6$. The case $r=3$ is not difficult to prove. Ramírez Alfonsín proved the case $r=4$ in 1999 [R99] and recently Hernández-Ortiz et al. proved it for $r=5$ [HOKMS23]. In 2015 Montejano and Ramírez Alfonsín proved the conjecture for Lawrence oriented matroids [MR15]. Furthermore, Bárány et al.[BBLP95] proved that for fixed $r$, the number of $1$-neighborly reorientations of any rank $r$ realizable oriented matroid on $n$ elements is $2\binom{n-1}{r-3}+2O(n^{r-4})$, i.e., Roudneff’s conjecture holds asymptotically in the realizable setting (here we also point out that the authors proved this result in terms of hyperplane arrangements in the Euclidean space and not in terms of realizable oriented matroids). In this manuscript, we prove that Roudneff’s conjecture holds asymptotically for all oriented matroids (Section 3, Corollary 3.10).
Conjecture 1.1 was stated for $n\geq 2r-1$ probably because in 1991 the value of $c_{r}(n,1)$ was unknown for n<2r1n<2r-1 but in 2001 was computed its value for $n\geq r+1$ [FR01]. So, in view of Roudneff’s conjecture, the authors asked in [MR15] if Roudneff’s conjectured holds for $n\geq r+1$ [MR15, Question 2] and again, it turns out that it is true for $r\leq 5$ [HOKMS23] and for Lawrence oriented matroids [MR15]. For an integer $k\geq 0$ it was proposed in [HKM24] the following strengthening of 1.1.
@@
Firstly, notice that $f_{\mathcal{M}}(r,n,k)=f_{\mathcal{M}^{\prime}}(r,n,k)$ if $\mathcal{M}^{\prime}\in[\mathcal{M}]$. So, it is enough to check Conjecture 1.2 for the reorientation classes $\mathcal{M}$] of $\mathcal{M}$. Secondly, every rank $r$ oriented matroid $\mathcal{M}$ on $n$ elements can be perturbed to become a uniform rank $r$ oriented matroid $\mathcal{M}^{\prime}$ on $n$ elements, see [BVSWZ99, Corollary 7.7.9] and in particular $f_{\mathcal{M}}(r,n,k)\leq f_{\mathcal{M}^{\prime}}(r,n,k)$ (see [HKM24, Theorem 3.16]). Thirdly, it is well-known that there is exactly one reorientation class of uniform rank $r$ oriented matroids on $n\leq r+2$ elements (see [HKM24, Remark 1.3] and [BVSWZ99, Section 6.1]). Thus, we obtain the following observation.
@@
For $k=1$ 1.2 combines Roudneff’s conjecture and [MR15, Question 2]. Hence, for $k=1$ the answer is positive if $r\leq 5$ and also holds for Lawrence oriented matroids. Conjecture 1.2 was also proved in [HKM24] for $k=2$ and $r=6$, and also for every odd rank $r=2k+1$ [HKM24].
One might think that as in the Upper Bound Theorem [Mac] the reorientation class of $\mathcal{C}_{r}(n)$ is unique in attaining the maximum in 1.2. However, there are different reorientation classes that attain $c_{r}(n,k)$, for $r=5,n=8,9$ and $k=2$ and also for $r=7,n=10$ and $k=3$ [HKM24, Theorem 4.1]. Nevertheless, for $k=1$ it was conjectured in [HOKMS23] that the reorientation class of ${C}_{r}(n)$ is the unique class with $c_{r}(n,k)$ $k$-neighborly reorientations.
@@ 2. $k$-neighborly reorientations of LOMs
In this section we study the parameter $f_{\mathcal{M}}(r,n,k)$ when $\mathcal{M}$ is a Lawrence oriented matroid (LOM). We obtain obtain some results on $k$-Roudneff’s conjecture for LOMs.
Before defining a LOM, we need to introduce the chirotope of an oriented matroid. The definition of the base axioms of an oriented matroid is due to Gutierrez Novoa [Gut65] who called the structure multiply ordered sets. Lawrence [L82] proved that this yields another axiom system for oriented matroids. The term chirotope of an oriented matroid is due to Dress [Dre85], who reinvented oriented matroids. An oriented matroid of rank $r$ is a pair $\mathcal{M}=(E,\chi)$ of a finite ground set $E$ and a chirotope $\chi:E^{r}\to\{+,0,-\}$ satisfying:
@@
If $\chi:E^{r}\rightarrow\{-,+\}$ is a chirotope, then $\mathcal{M}=(E,\chi)$ is a uniform oriented matroid. Moreover, if $E=\{1,\ldots,n\}$ and $\chi(B)=+$ for any ordered tuple $B=(b_{1}<\ldots<b_{r})$, then $\mathcal{M}=(E,\chi)$ is ${C}_{r}(n)$, the alternating oriented matroid of rank $r$ on $n$ elements. It is known that every oriented matroid $\mathcal{M}$ has exactly two basis orientations and these two basis orientations are opposite, $\chi$ and $\chi$ (see [BVSWZ99, Proposition 3.5.2]).